# D-Band 2D MIMO FMCW Radar System Design for Indoor Wireless Sensing

### v3

@@ II-C4 Accuracy Versus the Number of Array Elements
Figure 9: RMSE versus the number of radar array elements. The minimum RMSE achieved using the MUSIC and MVDR algorithms is presented.
III Hardware of 2D Radar Array
III-A Array Architecture
A 2D radar array with $N_{x}\times N_{y}$ virtual elements can be realized through different numbers of TX and RX elements, and their physical arrangements. For a large radar array implemented on chip and operating in the D-band, the array architecture has significant effects on the performance, including the losses and power consumption of the LO distribution network and the TX-to-RX leakage signals.
We consider the architecture shown in Fig. 10, which comprises two linear arrays of RX and TX elements. This architecture allows for a symmetric LO distribution to the TX and RX elements. This leads to correlated phase noise in the TX and RX signals, which suppresses phase noise in the RX [6]. The D-band antenna array can be implemented using the antenna-in-package (AiP) technology and directly flip-chip attached to the chip die [29]. This allows the integration of large radar arrays with high radiation efficiencies (over 80%). We use this architecture to estimate the loss of the LO distribution network, the required TX output power, and the power consumption of radar array.
III-B LO Signal Distribution
Generation and distribution of the LO signal to the RX and TX elements is a major challenge in the large radar arrays operating in mm-wave bands. The LO generation comprises a phase-locked loop (PLL) and frequency multiplier(s) with low spurs, wide bandwidth, good chirp linearity, and high DC-to-RF efficiency. The LO distribution network comprises the routing paths to the TX and RX elements, and the LO amplifiers to compensate for the losses and maintain the signal integrity. The reference signal can be generated by an external PLL. We consider all of these circuit blocks as the LO distribution network, as shown in Fig. 10.
Frequency multipliers can be used at different locations in the radar array to upconvert the low-frequency reference signal to the D-band LO signals. Generally, some of the frequency multipliers can be shared between the TX and RX paths, and some others can be used separately in each TX and RX system. We can envision two extreme cases for the deployment of frequency multipliers. In the first approach, the reference signal is unconverted using a single high-performance frequency multiplier (usually comprising multiple stages) to the final RF signals and is then distributed between the TX and RX systems. The drawback is the greater loss of the LO distribution network and the higher power consumption of the LO amplifiers. In the second extreme approach, the low-frequency reference signal is distributed between the TX and RX paths, and is upconverted by dedicated frequency multipliers in the all TX and RX systems. This can reduce the losses of the LO distribution network and the power consumption of the LO amplifiers. However, this needs a total of $N_{TX}+N_{RX}$ frequency multipliers which can lead to high power consumption for large radar arrays. The optimal architecture should be between these two extremes, but it depends on many factors on the circuit and process level. In this work, we use the first approach to perform system simulations (see Fig. 10).
We assume that the output power of the reference chirp signal after being upconverted by the frequency multiplier is $P_{\rm REF}$. The power of the reference signal is reduced by divisions in the paths and the passive losses, and reaches to $P_{\rm LO,TX}$ in the TX and $P_{\rm LO,RX}$ in the RX systems. Both of these LO power levels should be within certain ranges for their efficient operation. In the TX, the LO power should be high enough to push the PA into saturation and achieve high efficiency, as shown in Fig. 11(a). However, an excessively high LO power can degrade gain, power-added efficiency (PAE), defined as $\rm{PAE=(P_{out}-P_{in})/P_{DC}}$, and the linearity of the PA. In the RX, the LO power should be high enough to enable switching of the differential pair transistors operating as the mixer, as shown in Fig. 11(b). On the other hand, for excessively large LO power levels, the conversion gain of the mixer, defined as $\rm{CG=P_{out,IF}/P_{in,RF}}$, will degrade due to the operation of the transistor pair in the triode mode during most of the signal cycle [10]. Therefore, the LO power delivered to TX and RX should be maintained within optimal power ranges,
$P_{\rm LO,min}\leq P_{\rm LO}\leq P_{\rm LO,max},$ (19)
where the lower and upper limits of the LO power should be determined using properties of the circuit and fabrication process.
Figure 10: Physical architecture of the 2D MIMO radar array realized using two linear arrays of TX and RX elements. LO signal distribution network includes loss of the LO signal paths $L_{P}$, loss of the power dividers $L_{D}$, the amplifiers with power gain of $G_{\rm LO,TX}$ and $G_{\rm LO,RX}$, used for loss compensation.
Figure 11: Structure of typical mm-wave circuits used in the radar: (a) power amplifier circuit, and its efficiency $\eta$ and PAE versus input power, (b) mixer circuit and its conversion gain (CG) versus LO power, (c) amplifier circuit with power gain boosting, and the maximum power gain $G_{\rm max}$, unilateral power gain $U$, and the power gain efficiency $\mathbf{\gamma=G/{P_{\rm DC}}}$ of an amplifier stage versus power consumption.
The reference signal is highly attenuated by the divisions between paths and the loss of on-chip transmissions lines in the D-band. Therefore, the LO amplifiers are distributed along the routing paths to prevent the need for a single high-gain LO amplifier after the reference signal generator. We propose to use one LO amplifier in each path with a power gain of $G_{\rm LO,TX}$ in the TX paths and $G_{\rm LO,RX}$ in the RX paths (see Fig. 10). The loss of the LO paths is dependent on their physical length, layout structure, and properties of the chip fabrication process. We assume that the length of all branches is $\Delta l$ to simplify derivations. This length is dependent on the routing layout, but it can be fairly approximated by a fraction of the signal wavelength, for example, $\Delta l\approx\lambda/2$. The loss per unit length of the LO distribution paths can be estimated using the loss properties of transmission lines in the CMOS process (typically 1–3 dB/mm in the D-band). The loss of each path with the length of $\Delta l$ is derived as $L_{P}$.
We can derive the LO power delivered to each of the TX and RX systems using Fig. 10 as
${P_{\rm LO,TX}}=\frac{L_{D}(G_{\rm LO,TX}L_{P})^{\log_{2}{N_{TX}}}}{N_{TX}}P_{REF}$ (20)
${P_{\rm LO,RX}}=\frac{L_{D}(G_{\rm LO,RX}L_{P})^{\log_{2}{N_{RX}}}}{N_{RX}}P_{REF},$ (21)
We can set the gain of LO amplifiers to completely compensate for the losses or even boost power of the reference signal if required. However, a higher gain of the LO amplifiers operating in the D-band can lead to significantly increased power consumption of the radar array. Therefore, gain of the LO amplifiers should be optimized based the required LO power and power consumption budget.
The total number of LO amplifiers in the TX path can be calculated using Fig. 10 as
which using $\sum_{k=0}^{N}r^{k}=(r^{N+1}-1)/(r-1)$ is derived as
$N_{\rm LO,TX}=2N_{TX}-1.$ (25)
Similarly, for the RX path, the number of LO amplifiers can be derived as
$N_{\rm LO,RX}=2N_{RX}-1.$ (26)
These results will be used to estimate power consumption of the LO distribution network in Section III-D.
III-C TX Output Power
In the D-band, the efficiency of the silicon power amplifiers (PAs) is very low (typically 5–10%) [6, 8, 10, 11, 13, 30]. This can lead to huge power consumption for a large radar array. Therefore, it is essential to keep the required TX output power low using proper design strategies.
The required TX output power can be derived using (15), (16), and (17) as
This indicates that for $p<4$, the required TX output power increases with the distance (see Fig. 4). Conventional DOA estimation approaches assume that the SNR is independent of the target distance, that is $p=0$. However, (27) indicates that this requires the TX power to be increased proportional with $d^{4}$. This can result in high power consumption in long indoor distances. In our proposed approach, we scale both the SNR and the TX output power with distance by selecting $p=2$. This can significantly reduce the required TX output power and the power consumption of the radar array.
III-D Power Consumption
In the silicon-based radar arrays operating beyond 100 GHz, the chip power density is significantly higher than that of the lower mm-wave bands, as a result of more compact chip size and reduced antenna spacing. The low power efficiency and gain of transistors, and the high loss of passive elements in the D-band further increase the power consumption of the radar array. Therefore, it is essential to develop a strategy to save the power consumption.
The power consumption of the radar array is dependent on the number of array elements, system architecture of the radar TX and RX, LO distribution network, circuit structures, and fabrication process. In the D-band, power consumption of the RX circuits is comparable with that of the TX circuits, and the LO distribution network can have a significant contribution which dominates the total power consumption. The power consumption can be estimated as
$P_{\rm{DC}}\approx P_{\rm{DC,TX}}+P_{\rm{DC,RX}}+P_{\rm{DC,LO}}.$ (28)
III-D1 TX Power Consumption
$P_{\rm{DC,TX}}(d)\approx\frac{N_{\rm{TX}}P_{\rm{TX}}(d)}{\eta_{\rm{TX}}}.$ (29)
In the radar array architecture considered in this work, the TX efficiency can be approximated by the PA efficiency, which is in the order of 5–10% for silicon-based PAs in the D-band [6, 8, 10, 11, 13, 30]. It is assumed that the TX circuits are reconfigured as the TX output power is scaled with the distance so that its efficiency remains almost the same. For the FMCW radars where a saturated PA can be used, this is much less challenging than the communication systems using the amplitude-modulated signals. As a result, the power consumption of the TX system scales with the distance, proportional with the TX output power.
III-D2 RX Power Consumption
The power consumption of these circuits is dependent on many factors on the circuit and process level, and, as a result, cannot be accurately predicted before the circuits are designed. We can assume a fixed power consumption for each RX and scale it with the number of RX elements in the array. To find a practical estimate, we investigated the power consumption of the TX, RX, and LO parts in the D-band radars presented in the literature and noted that the power consumption of each RX element is usually on the order of 0.5x-1x the power consumption of each TX element [6, 9, 10, 12]. Therefore, we estimate the power consumption of an RX element using that of an TX element as $P_{\rm{DC,RX}}/N_{\rm{RX}}\approx\alpha P_{\rm{DC,TX}}/N_{\rm{TX}}$ with $\alpha\approx 0.75$. We should note that the RX power consumption is not scaled with the TX output power (unlike the TX). Therefore, $P_{\rm{DC,TX}}$ should be evaluated using (29) at a fixed target distance, such as the reference distance $d_{0}$.
III-D3 LO Power Consumption
The LO distribution network comprises the LO amplifiers and frequency multiplier(s). The LO amplifiers are designed to achieve the maximum power gain, which is different with the design criteria of the PA. We define an efficiency metric namely the power gain efficiency for an LO amplifier as ratio of the power gain to the power consumption
$\gamma=\frac{G}{P_{\rm DC}}.$ (31)
In mm-wave bands above 100 GHz, the gain of transistors is very low and usually an embedding passive network is used to boost it [31, 32]. The maximum power gain of transistor is given by
$G_{\rm max}=(\sqrt{U}+\sqrt{U-1})^{2},$ (32)
where $U$ is the unilateral power gain (Mason’s invariant) that depends on the size and bias of the transistor and frequency [31]. By increasing the bias current, $U$ initially increases but after a certain current level it is saturated. Therefore, we can assume that the transistor is biased at an optimum current with an associated unilateral power gain of $U_{opt}$. This also depends on the size of CMOS transistors. Using (31) and (32), we can conclude that this optimal current leads to the maximum $\gamma$ [see Fig. 11(c)]. Usually, multiple amplifier stages should be cascaded to achieve the desired power gain.
$P_{\rm DC,LO}=\frac{(2N_{TX}-1)G_{\rm LO,TX}}{\gamma_{\rm LO,TX}}+\frac{(2N_{RX}-1)G_{\rm LO,RX}}{\gamma_{\rm LO,RX}}.$ (33)
This equation along with (22) and (23) indicate that the power consumption of the LO network is dependent on the number of TX and RX elements, the power gain from the reference to the LO signals, that is $P_{\rm TX,LO}/P_{\rm REF}$ and $P_{\rm RX,LO}/P_{\rm REF}$, the losses of the passive components, $L_{P}$ and $L_{D}$, and the power gain efficiency of the transistors, $\gamma_{\rm LO,TX}$ and $\gamma_{\rm LO,RX}$.
We should also include the power consumption of the frequency multiplier(s), $P_{\rm DC,M}$, in (33) for more accurate results. As we have considered the array architecture with a single frequency multiplier, its power consumption is not scaled with the number of TX and RX elements.
III-D4 Simulations
The power consumption of the TX, RX, and LO distribution network versus the number of radar array elements $N_{\rm TX}N_{\rm RX}$ is shown in Fig. 12. The following assumptions are applied. The TX output power and efficiency are 10 dBm and 10%, the target distance is 1 m, and power consumption of the frequency multiplier is 50 mW. The LO distribution network provides 10 dB power gain from the reference signal to the LO signals, that is $P_{\rm LO,TX}/P_{\rm REF}=P_{\rm LO,RX}/P_{\rm REF}=10$ in (22) and (23). The losses of the LO paths and the power dividers $L_{P}=2\,{\rm dB}$ and $L_{D}={\rm 1\,dB}$. The power consumption of the LO distribution network is shown for two values of power gain efficiency. For example, $\gamma={\rm 0.4\,(mW)^{-1}}$ means that the LO amplifiers offer power gain of 4 (6 dB) for 10 mW power consumption.
Fig. 12 indicates that the power consumption of the LO distribution network is dominant when the number of array elements is small. In this condition, each LO amplifier should provide a high gain to meet the required gain from the reference signal to the LO signals. This increases their power consumption. By increasing the number of array elements, two competing effects should be considered. First, the required gain of each LO amplifier is lower, as more LO amplifiers are cascaded and higher gain is achieved. This effect tends to reduce the power consumption of the LO amplifiers. Second, as the number of LO amplifiers increases with the number of array elements, (25) and (26), this effect tends to increase the power consumption. These two competing trends, can also be noted in (33), initially reduce the LO power consumption, leading to a flat area shown in Fig. 12, and then increase the LO power consumption as the second effect dominates.
The power consumption of the TX and RX systems increase proportional with the square root of the number of array elements (assuming $N_{\rm TX}=N_{\rm RX}$). For large number of array elements, the power consumption is dominated by the TX power consumption if the LO amplifiers feature a sufficient power gain efficiency. Otherwise, the power consumption of the LO distribution network will be dominant.
Figure 12: Power consumption of the TX, RX, and LO distribution network versus the number of radar array elements. LO power consumption is shown for two values of power gain efficiency (PGE) $\gamma$.
IV Indoor Wireless Sensing Scenarios
We consider the free-space and through-wall indoor wireless sensing scenarios. It is assumed that there is a dominant line-of-sight (LOS) wave propagation path between the radar array and the target, and the multipath effects are negligible. In the mm-wave bands above 100 GHz, the multipath signals are severely attenuated by multiple reflections from the indoor objects and walls, and their effect is less significant than in the lower frequency bands [2].
In the following simulations, the radar parameters are assumed as $G_{TX}=G_{RX}=\rm{10\,dB}$, $\lambda=\rm{2.1\,mm}$ (at 140 GHz), $\sigma=\rm{100\,cm^{2}}$ as an estimate of the human hand RCS for gesture recognition applications [33], the chirp duration $T_{c}=\rm{10\,\mu s}$, and $\rm{NF_{RX}=10\,dB}$ (Table I). The radar frame comprises 100 chirps and the measurement time is $T_{meas}=100T_{c}={\rm{1\,ms}}$.
@@ IV Indoor Wireless Sensing Scenarios
Figure 13: Required output power per radar TX element versus SNR and distance in different wireless sensing scenarios: (a) free space, (b) clear glass, (c) drywall, (d) wood door. Loss of indoor materials in 140 GHz is extracted from the measurements (Table II) [2].
IV-A Free-Space Sensing
@@
In the free-space wireless sensing, the required TX output power versus the SNR and distance is shown in Fig. 13(a). The required TX output power can be read less than 0 dBm in most of the SNR and distance ranges. It reaches to a maximum of approximately 3 dBm at 30 dB SNR and 10 m distance. Comparing this with the TX output power of the state-of-the-art on-chip radars presented in Table I, indicates that the required TX output power in the free-space sensing is within the practical power levels of the available D-band radars. Furthermore, if we increase the SNR to 40 dB to improve the accuracy of the DOA estimation, or for a target with 10x smaller RCS, the required TX output power will increase to 13 dBm at the extreme distance of 10 m, which is still within the practical limits. It is important to note that these levels of required TX output power are achieved using our proposed scaling approach with distance. This investigation indicates the promise of high-resolution indoor wireless sensing in the free-space scenario.
IV-B Through-Wall Sensing
@@
The through-wall wireless sensing has been successfully realized in the low RF bands (below 8 GHz) [34]. In these bands, the wireless signals can pass through the indoor walls with a moderate loss, but the signal wavelength is too large ($\rm{\lambda>3\,cm}$) and leads to coarse resolutions. The D-band can provide very fine resolutions ($\rm{\lambda\approx 2\,mm}$), but its feasibility has not been explored in the literature yet. This inspires to use the developed model to investigate the required TX output power and the power consumption of the radar array for the through-wall wireless sensing in the D-band.
@@
The loss of three representative indoor materials measured at 140 GHz [2] is presented in Table II. We assume that the penetration loss through the material is dominant and neglect other losses, including the surface scattering and multipath propagation. The loss of wall medium should be applied twice in the radar SNR to account for the transmitted signal from the radar and the reflected signal from the target
${\rm{SNR_{\rm TW}}}={\rm{SNR}}\times L_{\rm{wall}}^{2},$ (34)
@@
TABLE II: Loss of indoor materials measured at 140 GHz [2].
@@
V Discussion
Figure 15: Power consumption of the radar array with 64 elements versus the RX SNR for different indoor materials. (a) 1 m distance, (b) 5 m distance.
A comparison of the required TX output power versus the RX SNR for different indoor materials is shown in Fig. 14. The results are shown at two distances 1 m and 5 m. Fig. 14(a) indicates that the required TX output power at the distance of 1 m is very low in the all conditions. This is a promising result for the future of large radar arrays in short indoor distances. However, the required TX output power at 5 m distance can become very high and impractical in some conditions, as shown in Fig. 14(b).
The power consumption of the radar array with 64 elements versus the RX SNR is shown in Fig. 15. In the low SNR, the requited TX power is low and the power consumption of the array is dominated by that of RX and LO, which are independent of the SNR. The required TX power becomes significant for the high SNR and large loss of indoor material, for example, the wood door in Fig. 15(a). This results in a power consumption sharply increasing with the SNR. The power consumption is higher for a longer target distance, as shown in Fig. 15(b).
We evaluate the results of our investigations with practical considerations on the algorithm and hardware levels.
V-A Number of Array Elements
The effective SNR of the MIMO radar can be improved by using more elements of the array to achieve a higher accuracy of DOA estimation. However, increasing the number of array elements can lead to several practical limitations, including higher power consumption, increased mutual coupling between antennas [24], self-interference, and spurs in the radar [23], greater loss in the LO signal distribution paths [35], and mismatch in the orthogonality of the MIMO channels. These challenges are more severe in the D-band, and must be mitigated through hardware and algorithm techniques.
V-B DOA Estimation
The evaluation of two popular DOA estimation algorithms, MUSIC and MVDR, indicates that the MUSIC outperforms the MVDR for large arrays, while the MVDR can provide superior performance for small arrays. In practice, the required computational power, memory, and simulation time are also important criteria in resource-constrained and real-time applications. Furthermore, the emerging deep learning based DOA estimation approaches [36] can provide higher performance and adaptation to operating conditions.
V-C Target RCS
In the simulations presented, the target RCS was assumed as $\sigma=\rm{100\,cm^{2}}$, as an estimate of the human hand RCS for gesture recognition applications [33]. In practice, the target RCS can exhibit significant variations due to the diversity of targets (for example, human hand for gesture sensing or human chest for vital signs monitoring) and their volatile orientation with respect to the radar array. This can result in large changes in the RX SNR [see (15)] and the TX output power requirements. A practical approach is to design the system for a range of target RCS values.
V-D Radar Measurement Time
The measurement time, $T_{meas}$ in (15), can be increased to achieve a higher RX SNR and lower the required TX output power (for example, a gain of 10 dB by increasing it from 1 ms to 10 ms). However, this results in a higher power consumption of the radar, more computational resources to process the received signal samples, and longer processing times. Therefore, the measurement time should be selected based on all of these criteria.
V-E TX Output Power Requirement
The TX output power is linearly dependent on the RX SNR through (15). The gain of TX and RX antennas as well as NF of the RX are usually limited by the implementation technology. The radar measurement time can be changed within the practical limits. In the through-wall sensing scenario, a higher TX output power is required to compensate for the extra losses. In practice, there are other losses due to wave scattering by the material in the D-band [2]. Therefore, the radar should be able to provide a wide range of TX output power.
V-F Power Consumption
The power consumption of the radar array is dominated by the RX and LO circuits under the condition that a low SNR is sufficient for the DOA estimation. However, in the high-SNR regime, the power consumption is significantly dependent on the SNR. In the large radar arrays implemented on silicon processes, the density of power consumption (per unit area) is very high in the D-band, and should be properly managed through the packaging and heatsink considerations.

### v4

@@ IV-B3 Accuracy of Adaptive Radar
Figure 9: RMSE versus the number of radar array elements. The minimum RMSE achieved using the MUSIC and MVDR algorithms is shown, i.e., $\rm{RMSE}=\min(RMSE_{MUSIC},RMSE_{MVDR})$.
IV-C Computational Efficiency
The power consumption of the MIMO radar back-end is dependent on the number of array elements, the size of collected data, the computational complexity of the algorithms, and the features of the digital signal processor (DSP). As most of these information are unknown, we use the relative simulation time as a metric for computational efficiency of the radar back-end. Simulations are performed using an Intel Core i5-7300HQ 2.50 GHz CPU with 8 GB RAM. The simulation time comprises the time required to complete the DOA estimation simulations for SNR of 0–30 dB with a step of 5 dB, distance of 1–10 m with a step of 1 m, and 100 Monte Carlo runs.
Simulation times for various radar array sizes and DOA estimation algorithms are shown in Fig. 10. For an $8\times 8$ array, the MUSIC algorithm provides much faster simulations, by a factor of about 2x, than the MVDR algorithm. The accuracy is also much higher for the MUSIC algorithm (see Fig. 7). Therefore, for an $8\times 8$ array the MUSIC algorithm provides the highest performance and energy efficiency. On the other hand, for a $2\times 2$ array, the MVDR algorithm outperforms the MUSIC algorithm in terms of the simulation time and also the accuracy. It can be concluded that for large arrays, e.g., over 16 elements, the MUSIC algorithm provides the highest performance while for small arrays, e.g., less than 16 elements, the MVDR algorithm prevails. The adaptive radar architecture allows dynamic reconfiguration of the array size and the selection of the DOA estimation algorithm.
Figure 10: Simulation times of the DOA estimation for various radar array sizes and algorithms.
@@ IV-C Computational Efficiency
Figure 11: Required TX output power (per radar element) versus the SNR and target distance in various wireless sensing scenarios at the D-band (140 GHz): (a) free space, (b) clear glass, (c) drywall, (d) wood door.
V Performance of Adaptive MIMO Radar for D-Band Wireless Sensing
We evaluate performance of the adaptive MIMO radar for D-band wireless sensing and compare it with a conventional MIMO radar. We consider the free-space and through-wall indoor wireless sensing scenarios. It is assumed that there is a dominant line-of-sight (LOS) wave propagation path between the MIMO radar and the target, and the multipath effects are negligible. In the mm-wave bands above 100 GHz, the multipath signals are severely attenuated by multiple reflections from the indoor objects and walls, and, as a result, their effect is less important than in the lower frequencies [2].
In the following simulations, the radar parameters are assumed as $G_{TX}=G_{RX}=\rm{10\,dB}$, $\lambda=\rm{2.1\,mm}$ (at 140 GHz), $\sigma=\rm{100\,cm^{2}}$ as an estimate of the human hand RCS for gesture recognition applications [33], the chirp duration $T_{c}=\rm{10\,\mu s}$, and $\rm{NF_{RX}=10\,dB}$. The radar frame comprises 100 chirps and the measurement time is $T_{meas}=100T_{c}={\rm{1\,ms}}$. The TX output power and the RX SNR are scaled with the distance based on a scaling exponent of $p=2$.
Figure 13: Power consumption of the MIMO radar with 64 elements versus the RX SNR for different indoor materials. (a) 1 m distance, (b) 5 m distance.
V-A Free-Space Sensing
@@
In the free-space wireless sensing, the required TX output power versus the SNR and distance is shown in Fig. 11(a). The required TX output power can be read less than 0 dBm in most of the SNR and distance ranges. It reaches to a maximum of approximately 3 dBm at 30 dB SNR and 10 m distance. The required TX output power in the free-space sensing is within the practical power levels of the available D-band radars. Furthermore, if we increase the SNR to 40 dB to improve the accuracy of the DOA estimation, or for a target with 10x smaller RCS, the required TX output power will increase to 13 dBm at the extreme distance of 10 m, which is still within the practical limits. It is important to note that these levels of required TX output power are achieved using our proposed scaling approach with distance. This investigation indicates the promise of high-resolution indoor wireless sensing in the free-space scenario.
V-B Through-Wall Sensing
@@
The through-wall wireless sensing has been successfully realized in the low RF bands (e.g., sub-6 GHz) [34]. In these bands, the wireless signals can pass through the indoor walls with a moderate loss, but the signal wavelength is too large ($\rm{\lambda>3\,cm}$) and leads to coarse resolutions. The D-band can provide very fine resolutions ($\rm{\lambda\approx 2\,mm}$), but its feasibility has not been explored in the literature yet. This inspires to use the developed model to investigate the required TX output power and the power consumption of the MIMO radar for the through-wall wireless sensing in the D-band.
@@
The loss of three representative indoor materials measured at 140 GHz [2] is presented in Table I. We assume that the penetration loss through the material is dominant and neglect other losses, including the surface scattering and multipath propagation. The loss of wall medium should be applied twice in the radar SNR to account for the transmitted signal from the radar and the reflected signal from the target
${\rm{SNR_{\rm TW}}}=\frac{\rm{SNR}}{L_{\rm{wall}}^{2}},$ (32)
@@
TABLE I: Loss of indoor materials measured at 140 GHz [2].
@@
A comparison of the required TX output power versus the RX SNR for different indoor materials is shown in Fig. 12. The results are shown at two distances 1 m and 5 m. The required TX output power at the distance of 1 m is very low in the all conditions. This is a promising result for the future of large MIMO radars in short indoor distances. However, the required TX output power at 5 m distance can become very high and impractical in some conditions.
The power consumption of the MIMO radar with 64 elements versus the RX SNR is shown in Fig. 13. In the low SNR, the requited TX power is low and the power consumption of the array is dominated by that of RX and LO, which are independent of the SNR. The required TX power becomes significant for the high SNR and large loss of indoor material, e.g., the wood door. This results in a power consumption sharply increasing with the SNR. The power consumption is higher for a longer target distance.